Suppose you invest $1,050 in an account paying 8% interest per year a. What is the balance in the account after 2 years? How much of this balance corresponds to interest on interest? b. What is the balance in the account after 26 years? How much of this balance corresponds to interest on interest?
Answer #1

Balance in account after 2 years = 1.050(1.08)2

= 1,224.72

Balance on account if the interest was not compounded = 1050 + 1050*8%*2 = $1,218

Interest on Interest = $1,224.72 – 1,218

= $6.72

b.Balance in Account after 26 years = 1.050(1.08)26

= $7,766.17

Balance on account if the interest was not compounded = 1050 + 1050*8%*26

= $3,234

Interest on Interest = $7,766.17 - $3,234

= $4,532.17